OpenStudy (anonymous):

This next question is a little different. Write the equation in slope-intercept form of the line that has a slope of 3 and contains the point (4, 9).

OpenStudy (anonymous):

y-y1 = m(x-x1) x1,y1 = 4,9 y-9 = 3(x-4) y-9 = 3x-12 y = 3x-3

OpenStudy (anonymous):

Thanks(:

OpenStudy (anonymous):

ye